OpenStudy (anonymous):

For the investment with interest compounded annually, find the final balance using the formula A = P(1 + r)t. $600 for 4 years at 5% A. $120 B. $129.30 C. $720 D. $729.30

OpenStudy (1645323):

720

OpenStudy (texaschic101):

A = P(1 + r) t P = 600 r = 0.05 t = 4 now we sub A = 600(1 + 0.05)^4 A = 600(1.05^4) A = 600 (1.2) A = 720 1645323 is correct....answer is 720